The mass percent of oxygen in 109% oleum is???
A-40%
B-63.18%
C-60%
D-24%

Grade:11

1 Answers

Vikas TU
7 years ago
 109% oleum  means we have now 109 gm of H2SO4.
Therfore, moles of H2SO4  = 109/98 = 1.112 moles
moles of O atoms = 1.112*4 = 4.448 moles.
weight = 4.448*16 = 71.18 gm.
% = 71.18/109 * 100
= 65.30 %